# Human Reactions to Incorrect Answers from Robots

### v1

@@ V Results and Discussion
To provide robust experimental evidence, a one-way ANOVA test [13] was conducted to analyze the change in participants’ trust levels before and after the experiment. The results, presented in Fig 6, demonstrate a highly significant difference in mean trust levels. In this visualization, we observe that prior to our experiment, the peak density of human trust in the robot hovered around 0.5, but post-experiment, it noticeably increased to around 1.5. The one-way ANOVA test reveals a significant overall change in trust levels across experimental groups, supported by a high F-statistic of 150.69 and a corresponding p-value of $3.16\times e^{-13}$.
@@
Completely trusted (a)
7
0
0
Moderately trusted (b)
0
11
0
Slightly trusted (c)
0
0
1
To compare the trust levels before and after the experiment, we created a contingency table. Table I allows us to visualize the relationship between the trust level before the experiment (rows) and the trust/reliability level after the experiment (columns). Based on the table, we can confidently say that the trust level increases after the experiment, indicating that the participants perceived the robot as more reliable following the experimental intervention.
@@
According to the chi-square statistic it is clear that there is a significant association between the trust levels before and after the experiment.
This research provide the following additional insights from our participants into specific aspects of human-robot interactions:
V-1 Trust Impact of Unacknowledged Errors
Data indicates that 77% of participants experienced a decrease in trust when errors were not acknowledged. This finding underscores the importance of transparent communication in maintaining user trust in robotic systems.
V-2 Importance of Acknowledging Errors
All of the participants emphasized the significance of robots admitting mistakes which means acknowledging errors appears to play a crucial role in maintaining user trust.
V-3 Effeteness of Robot’s Apology
The 87% of participants expressed that after making a mistake, robot’s apology is very effective to gain the trust.
V-4 Error Impact on Trust
Our analysis suggests that, for 62% of the participants, the errors have a negative impact on trust. However, the extent of this impact varies among participants, with some individuals displaying resilience to errors.
V-5 Perceived Reliability of Robots
Experimental result suggests that, after the interaction with the robot 84% of participants perceive robots as reliable, aligning with the initial expectation of participants finding robots somewhat reliable and useful.
V-6 Future Interaction Intent
We investigated the participants’ intent for future interactions with robots. 93% of participants expressed willingness to engage with robots in future tasks, indicating a positive outlook on the integration of robots into various domains.
@@ VI Conclusion and Future Work
In this study, we found that participants’ degrees of confidence in robots varied, with many exhibiting some degree of trust even in the case of those who had not before interacted with robots. The experimental results support both of our hypothesises (H1 and H2). The majority of the participants were aware of the robot’s shortcomings, and their level of trust in it was influenced by whether or not the robot admitted its mistakes. Even with fault knowledge, there was a noticeable willingness to interact with robots in subsequent assignments. Even when they made mistakes, participants tended to believe that robots were either equally or even more trustworthy, indicating that public acceptance of robots is growing. The study demonstrates that people’s reactions to interacting with robots are nuanced, weighing doubt against an appreciation of the potential benefits of robotics. In our upcoming research, we will investigate how people’s long-term trust and dependence on robots are impacted by the kind and frequency of errors made by them. We’ll also investigate how people from various demographic groups and cultures respond to engaging with robots. Lastly, to increase robot dependability and user trust, we will strive to make them even more adept at identifying and correcting mistakes. In future, we are planning to do the same experiment on different types of robots to see how people react on different types of robot and how different types of robots effect on human trust on automation.
@@ ACKNOWLEDGMENTS
We appreciate all of the help that was provided by Dr. David Feil-Seifer and Dr. Monica Nicolescu for sharing their opinions with us. We also appreciate our participants those who help us by participating in our survey.

### v2

@@ V-C Statistical Analysis of Trust Level Changes
To provide robust experimental evidence, we conducted a one-way ANOVA test [13] to analyze the change in participants’ trust levels before and after the experiment. The results, visualized in Figure 8, reveal a highly significant increase in mean trust levels post-experiment.
As seen in Figure 8, the density distribution before the experiment (blue) shows a broader spread with a lower peak, indicating a more varied and generally lower level of trust in the robot. In contrast, post-experiment (red), the distribution becomes narrower and sharply centered around a higher trust level, suggesting a more consistent and elevated perception of reliability.
The one-way ANOVA test confirms a statistically significant overall change in trust levels across experimental groups, yielding an F-statistic of 150.69 and a p-value of 3.16 × 10-̂13. This highly significant result indicates that participant trust in the robot significantly increased after the experiment.
@@
Completely trusted (a) 7 0 0
Moderately trusted (b) 0 11 0
Slightly trusted (c) 0 0 1
To compare the trust levels before and after the experiment, we created a contingency table. Table I allows us to visualize the relationship between the trust level before the experiment (rows) and the trust/reliability level after the experiment (columns). Based on the table, we can confidently say that the trust level increased after the experiment, indicating that the participants perceived the robot as more reliable following the experimental intervention.
@@
The chi-square analysis ($\chi^{2}=38.0$, $p=1.12\times 10^{-7}$) confirms a significant association between pre-experiment and post-experiment trust levels. This result validates that the observed increase in trust was not random but statistically meaningful, reinforcing the effectiveness of error acknowledgment in human-robot trust building.
V-D User Perspectives on Trust and Errors
This research provides additional insights into key aspects of human-robot interaction based on participant responses.
V-D1 Trust Impact of Unacknowledged Errors
This research indicates that 77% of participants experienced a decrease in trust when errors were not acknowledged. This finding underscores the importance of transparent communication in maintaining user trust in robotic systems. When a robot failed to acknowledge its mistakes, participants perceived it as less reliable, reinforcing the idea that users expect robots to be accountable for their errors.
V-D2 Importance of Acknowledging Errors
All participants emphasized that robots admitting mistakes is crucial for trust maintenance. Acknowledgment of errors not only enhances transparency but also reassures users that the system is aware of its limitations. These findings suggest that designing robots with socially appropriate error-handling mechanisms can improve their perceived reliability and overall trustworthiness.
V-D3 Effectiveness of Robot’s Apology
87% of participants found that a robot’s apology after making a mistake was effective in restoring trust. This result suggests that apologetic behaviors can mitigate negative perceptions of robotic errors, making robots more acceptable in real-world applications. The ability of robots to acknowledge mistakes and issue apologies helps foster a more human-like interaction, reducing skepticism and reinforcing user confidence.
V-D4 Error Impact on Trust
Errors had a notable impact on trust, with 62% of participants reporting a negative effect on their perception of the robot’s reliability. However, this impact was not uniform across all participants. While some individuals were more resilient to errors, others became more skeptical of the robot’s capabilities. This variation suggests that users’ prior experiences and expectations of automation influence how they perceive robotic reliability.
V-D5 Perceived Reliability of Robots
Despite concerns about errors, 84% of participants regarded robots as reliable after interacting with them. This aligns with the initial expectation that robots, though imperfect, are still functional and useful. The findings suggest that robot reliability is judged not just by error occurrence but also by how the robot handles mistakes.
V-D6 Future Interaction Intent
93% of participants expressed willingness to engage with robots in future tasks, demonstrating a positive outlook on human-robot interaction. Despite encountering errors, participants remained open to using robotic systems, suggesting that trust in automation can be strengthened through effective error-handling strategies and transparent communication. The findings highlight the importance of designing robots that not only perform tasks efficiently but also manage errors in a human-like and socially acceptable manner.
@@ VI Conclusion and Future Work
In this study, we found that participants’ confidence in robots varied, with many exhibiting some degree of trust, even among those who had never interacted with robots before. Our experimental results support both hypotheses (H1 and H2).
Regarding H1, we observed that trust in robots was influenced by whether the robot acknowledged its mistakes. The majority of participants recognized the robot’s limitations, and 84% reported increased trust when errors were admitted and apologized for. Despite being aware of errors, participants remained willing to interact with robots in future tasks, highlighting the importance of transparent error acknowledgment in trust-building.
For H2, our results confirm that humans are generally more forgiving of errors made by other humans or search engines than those made by robots. 55% of participants were less forgiving of robotic mistakes, suggesting that people hold robots to higher accuracy standards than they do humans. However, when robots admitted their mistakes, trust improved, indicating that error-handling mechanisms play a crucial role in shaping user perceptions of robotic reliability.
Interestingly, even when mistakes were made, participants often perceived robots as equally or more trustworthy, suggesting that public acceptance of robots is increasing. Our findings demonstrate that people’s trust in robots is nuanced, balancing skepticism with an appreciation of their potential benefits.
For future research, we aim to investigate how long-term trust and reliance on robots are influenced by the type and frequency of errors they make. We will also examine how individuals from diverse demographic and cultural backgrounds respond to robot interactions. Additionally, to improve robot dependability and user trust, we will work on enhancing their ability to recognize and correct mistakes more effectively.
Furthermore, we plan to conduct similar experiments with different types of robots to analyze how people react to various robotic designs and to assess how different robot types influence human trust in automation. These studies will provide deeper insights into the social and cognitive factors shaping trust in robotic systems.
@@ ACKNOWLEDGMENTS
We appreciate all of the help that was provided by Dr. David Feil-Seifer for sharing their opinions with us. We also appreciate the participants who helped us by participating in our survey.